OpenStudy (anonymous):

Determine whether ln(n/(3n+1)) diverges or converges.

OpenStudy (anonymous):

*

OpenStudy (anonymous):

diverges or converges where? at x=0? at x=∞?

OpenStudy (anonymous):

as x--> infinity

OpenStudy (anonymous):

I'm pretty sure that converges, but I forget how to explain why.

OpenStudy (turingtest):

\[\lim_{x\to a}\ln(f(x))=\ln(\lim_{x\to a}f(x))\]

OpenStudy (turingtest):

can you take\[\lim_{n\to\infty}\frac n{3n+1}\]?

OpenStudy (anonymous):

Wouldn't that be indeterminate, infinity/infinity?

OpenStudy (turingtest):

not if you divide numerator and denominator by n first

OpenStudy (anonymous):

The +1 in the denominator doesn't make a difference at infinity, so take n/(3n).

OpenStudy (turingtest):

^another valid way to see it

OpenStudy (anonymous):

I sometimes forget how to run the numbers, but I can visualize my way through most anything! ;-)

OpenStudy (anonymous):

ohh ok, I see. So that would allow me to see that the limit is 1/3. I am supposed to use a convergence test to show this though.

OpenStudy (anonymous):

If you want to see where the limit converges to, take LN(1/3).

OpenStudy (turingtest):

yes, this *is* the convergence test; checking if the limit converges

OpenStudy (anonymous):

Oh, I have a list from my prof of the root test, ratio test, comparison test, p-series test, integral test, geometric series test, and test for divergence. I suppose though if I can find the limit that is proof that the series converges

OpenStudy (turingtest):

oh you are doing a series!?

OpenStudy (turingtest):

then the series only converges if the limit of it's terms is 0, which here it clearly is not

OpenStudy (anonymous):

So as a series it diverges?

OpenStudy (turingtest):

yes, think about what it means for the terms of the series to converge to a finite number as we approach infinity this would be like adding \[\ln\frac13+\ln\frac13+\ln\frac13+\ln\frac13+....\]for infinity. this will obviously not converge as we keep adding the same number to it

OpenStudy (anonymous):

Ah, that explains the 'n' instead of 'x'

OpenStudy (turingtest):

if the limit of the terms of the series converge to zero the series will end up adding \[0+0+0+0+...\]as \(n\to\infty\), so the series can be finite as we don't keep adding stuff to it make sense?

OpenStudy (anonymous):

It just clicked. I understand now, using the nth term test for divergence I can justify my reasoning. Thanks!

OpenStudy (turingtest):

Happy to help :)